OpenStudy (anonymous):

how do you solve g(a)=a^2-3a, f(a)=2a+4 Find g(-7) +f(-7). Composition Functions.

OpenStudy (anonymous):

=g(-7)+f(-7) ={(-7)^2-3*(-7)}+{2*(-7)+4} =49+21-14+4 =60

OpenStudy (anonymous):

thats it!

OpenStudy (anonymous):

thanks!